OpenStudy (anonymous):

The length of the top of a table is 4m greater than the width. The area is 60m^2. Find the dimensions of the table.

OpenStudy (anonymous):

Let x m be the length of the table then width of the table is (x-4)m, Now, Area of rectangle = length * width 60 =x * (x-4)

OpenStudy (anonymous):

Now, solve for x can u?

OpenStudy (anonymous):

x(x - 4) = 60 x^2 - 4x - 60 (x -10) (x + 6) = 0 x = 10 and -6?